# An ocean front detection and tracking algorithm

### v1

@@ Abstract
Ocean front is defined as the interface between different water masses and plays a vital role in the evolution of many physical phenomena. Previous detection methods are based on histogram, Lyapunov exponent, gradient and machine learning. These algorithms, however, introduce discontinuity, inaccuracy, use less information or just approaching traditional results. Moreover, automatic front tracking algrorithm is not open source in preceding studies. This paper foucuses on large-scale ocean fronts and proposes an automatic front detection and tracking algorithm based on Bayesian decision and metric space. In this, front merging, filling and ring deletion are put forward to enhance continuity. The distance between fronts in different days is firstly defined and is well-defined in metric space for functional analysis. These technologies can be migrated to other areas of computer vision such as edge detection and tracking.
@@ I Introduction
Ocean front is defined as the interface of water masses with different properties such as temperature, salinity, density and chlorophyll[1][2]. With the fast development of oceanic observation and computer vision, ocean front detection and tracking techniques, no matter manual or automatic, have been uplifted to an unprecedented level.
There are mainly four genres of ocean front detection algorithm: histogram, Lyapunov, gradient and machine learning[3]. The histogram-based algorithm is based on classical variance analysis used in a window of some property, i.e., the ratio of between-variance and within-variance is calculated and a F-test is executed[4][5].
For every attempted value between mininum and maximum value in the window, a F-score is calculated and the maximum F-score is compared with a threshold[6]. This algorithm has been improved through cross-validation but still suffer the imputation of not handling multi-modal patterns well and over-detection[7].
Apart from Eulerian perspectives, the Lagrangian indicators such as finite time Lyapunov exponent(FTLE) and finite size Lyapunov exponent(FSLE) can also detect fronts[8][9]. If two manifolds(water masses) generate around the hyperbolic fixed point(HFP), then they will follow the streamline to spread out without crossing over each other[10][11]. A front is thus formed between the manifolds and reasearchers note the place where HFP occurs as a front. HFP and elliptical fixed point(EFP) are jointly generated, so the front found is around an eddy, i.e, this method can only determine a subset of true fronts[12][13][14].
Though noted for its pithiness and highly precise, supervised learning algorithm bears the blame for merely approximating the other algorithms[15][16][17][18]. Without using other methods’ results as labels, the unsupervised algorithm seems reasonable whereas the problem of setting the number of clusters exists[19].
Nearly all of the gradient method is based on Canny algorithm and is widely used by researchers to determine front[20]. There exists the problem of discontinuity for the information used is only gradient, i.e. fisrt-order derivative[21][22]. What is more, the results can be grotesque–rings for example[23][24].
With regard to front tracking, no one gives a clear definition and algorithm.
Therefore, the initial research object is changed in this paper and is based on gradient algorithm. It focuses on frontal zone, which is a broader range of front candidates rather than the narrow front with one pixel width. It seems reasonable for that water masses meet at a wider scale and the fronts change daily. Moreover, only the gradient information, i.e., the first order derivative is used in preceding researches. Here, the gradient is seen as a prior probability, it is combined with two opeartors calculated by the original field information to give a Bayesian decision[25][26][27]. This procedure eliminates the problem of setting a threshold mentioned above, thus avoiding the threshold-sensitive situations. After the frontal zones are detected, mathematical morphology is used to thin the front zones to fronts and redundant branches are removed[28][29]. In addition, deep first search algorithm is applied to follow every front and a merging method is developed. A ring-deletion algorithm is proposed to delete the ring structures iteratively to ensure the rationality. Three distances are defined in a metric space and applied to fronts in the former and latter day. Finally, distance-width decision is imposed to identify the same front in time series.
@@ II-D Bayesian decision
Secondly, likelihood is obtained by using more information of the whole field, i.e., local degree of edge(LDE) and block deviation(BD). LDE and BD are calculated for the $4$ pairs of points($A-I$, $B-H$, $C-G$ and $D-F$) around $E$.
@@
In (5), $V=[A,B,C,D,F,G,H,I]$ is the vector of $8$ SST values around $E$. $V_{max}$ is the maximum of these components in $V$, $V_{min}$ is the minimum and $\overline{V}$ the average. It can be verified that LDE and BD are two mathematical operators which fall into $[0,1]$ for normalization.
@@ II-E Skeletonization
$\begin{gathered}S(Z)=\bigcup_{k=0}^{K}S_{k}(Z)\\ S_{k}(Z)=(Z\ominus kM)-(Z\ominus k)\circ M\\ K=\max\{k|(Z\ominus kM)\neq\varnothing\}\end{gathered}$ (10)
The principle can be seen in figure 3, and the mathematical principle can be seen in (10). In it, $Z$ is the determined frontal zone, $(Z\ominus kM)$ means performing $k$ rounds of corrosion on the identified frontal zone, $M$ is the convolutional kernel, $\circ$ is the open operation and $K$ is the maximum number of corrosion cycles before being corroded into an empty set.
@@ II-F Skeleton trimming
Although these fronts have been found, they have many branches, and in order to extract the main features and adapt to the research scale, redundant branches must be removed. Here, a greedy algorithm called the discrete skeleton evolution(DSE) algorithm is used to assign weights to each branch, iteratively removing the branch with the minimum weight to obtain the main fronts.
@@ IV-B1 Detection
The higher occurrence of SST frontal zones in the northern parts of SCS, particularly in the latitudes between $10$°N and $20$°N, reflects seasonal transitions, where the differences between warm and cold ocean currents generate frontal zones. The seasonal shift in atmospheric circulation, coupled with oceanic upwelling and the influence of wind patterns, plays a crucial role in the formation of these frontal zones. The convergence of different water masses, such as warm tropical waters and cold subpolar currents, is responsible for the intensification of these frontal zones.
@@
Winter shows the highest intensity ($0.25$°C/km), followed by spring ($0.20$°C/km) summer ($0.15$°C/km) and fall ($0.14$°C/km). Intensity refers to the temperature gradient across the ocean front, which is typically stronger in spring and winter. Spring has more distinct temperature differences due to seasonal transitions, while winter’s cold temperatures can create higher gradients. The lower intensity values in summer and fall indicate less temperature contrast between the ocean and surrounding atmosphere.
@@
Figure 14: Daily average offshore distance of SST fronts from 2022 to 2024
Figure 15: Seasonal mean and standard deviation of offshore distance from 2022 to 2024
TABLE II: YEARLY AVERAGE OF OFFSHORE DISTANCE
Year 2022 2023 2024
Distance(km) 109.56 98.88 103.53
@@
Figure 14 and 15 show the daily and seasonal length-weighted average offshore distance of SST fronts respectively. The data exhibits pronounced seasonal oscillations, with offshore distance increasing during cooler seasons($109.30$km, $107.51$km for fall, winter)and decreasing in warmer seasons($97.99$km, $101.30$km for spring, summer). The standard deviation, however, differs greatly($13.37$, $15.27$, $16.85$ and $13.35$). The peak in fall is caused by more frequent weather changes, and the trough in winter and spring is attributed to relative stable weather conditions.
The drastic decline of offshore distance in 2023, as seen in table II, is consistent with the last El Nino phenomenon. During El Nino, the ocean’s thermal structure is changed, causing fronts closer to the shore. In addition, the induced regional climatic anomalies such as storms and wind shifts can also push SST fronts closer to the shore.
@@ IV-C Comparison
TABLE III: QUANTITATIVE COMPARISONS ON SST DATASET
@@
Table III presents a quantitative comparison of the four methods, the value is the average of detection results from 2022 to 2024. As to number, improved method significantly outperforms the other methods, detecting only $71.44$ fronts, which indicates a more accurate and less over-detecting performance. In contrast, histogram method detects $263.22$ fronts, indicating that it tends to over-detect front, leading to potential false results. Lyapunov and gradient method also detect more fronts than improved method, but still show better accuracy compared to histogram method.
Improved method yields the highest intensity of $0.16$°C/km, indicating a sharper detection of the fronts. Gradient method comes close with $0.093$°C/km, but both histogram and Lyapunov method have much lower intensity values, suggesting weaker and less defined front detection. This suggests that improved method is more capable of detecting stronger fronts, while histogram and Lyapunov methods tend to miss or under-represent the intensity of these fronts.
@@ V Conclusion
Quantitative comparisons demonstrate the superiority of the proposed method: it reduces over-detection by $73\%$ compared to histogram-based approaches, achieves sharper intensity ($0.16$°C/km), and captures broader ($40.56$ km) and longer ($212.42$km) fronts. These improvements stem from the fusion of multi-scale information and iterative refinement strategies, which traditional gradient or Lagrangian methods lack. Additionally, the open-source release of the tracking algorithm fills a critical gap in reproducible research.

### v3

@@ Abstract
Existing ocean front detection methods—including histogram-based variance analysis, Lyapunov exponent, gradient thresholding, and machine learning—suffer from critical limitations: discontinuous outputs, over-detection, reliance on single-threshold decisions, and lack of open-source implementations. To address these challenges, this paper proposes the Bayesian Front Detection and Tracking framework with Metric Space Analysis (BFDT-MSA). The framework introduces three innovations: (1) a Bayesian decision mechanism that integrates gradient priors and field operators to eliminate manual threshold sensitivity; (2) morphological refinement algorithms for merging fragmented fronts, deleting spurious rings, and thinning frontal zones to pixel-level accuracy; and (3) a novel metric space definition for temporal front tracking, enabling systematic analysis of front evolution. Validated on global SST data (2022–2024), BFDT-MSA reduces over-detection by $73\%$ compared to histogram-based methods while achieving superior intensity ($0.16\degree$C/km), continuity, and spatiotemporal coherence. The algorithm’s open-source release bridges a critical gap in reproducible oceanographic research.
@@ I Introduction
Ocean fronts, interfaces between distinct water masses, are pivotal in regulating marine ecosystems, heat transport, and extreme weather[1, 2]. Despite advances in remote sensing, existing detection methods exhibit fundamental flaws[3]. Histogram-based approaches over-detect fragmented fronts due to rigid windowed F-tests, while Lagrangian methods (e.g., Lyapunov exponents) capture only eddy-associated fronts, ignoring wind- or topography-driven features[4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Gradient thresholding, though widely adopted, produces discontinuous outputs and artificial rings, and machine learning merely replicates traditional results without innovation[15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Critically, no open-source framework exists for reproducible front tracking[3].
This paper addresses these gaps through three key contributions:
1. Unified Bayesian Detection: By redefining fronts as probabilistic zones rather than binary edges, manual thresholds are replaced with a Bayesian fusion of gradient priors and field operators, enhancing robustness to noise and multi-scale variability.
2. Morphological Coherence Enforcement: A suite of algorithms—including ring deletion, branch trimming, and zone thinning—eliminates geometric artifacts while preserving topological continuity, a first in ocean front studies.
3. Metric Space Tracking: Front evolution analysis is formalized by defining temporal distances in a metric space, enabling rigorous tracking of front identity and dynamics over time.
The proposed BFDT-MSA framework is validated against four baseline methods: number, intensity, width and length, demonstrating superior performance in reducing over-detection ($71$ vs. $263$ fronts), sharpening intensity ($0.16\degree$C/km), and capturing realistic front dimensions . Its open-source release ensures transparency and adaptability for diverse applications, from climate modeling to fisheries management.
Key advantages over prior work
Unlike threshold-dependent methods, BFDT-MSA leverages probabilistic reasoning to handle ambiguous gradients. Unlike Lagrangian approaches, it detects all front types—not just eddy-associated features. Unlike black-box machine learning, it provides interpretable, physics-guided outputs. These advancements establish a new paradigm for ocean front analysis.
@@ II-D Bayesian decision
Secondly, likelihood is obtained by using more information of the whole field, i.e., local degree of edge(LDE) and block deviation(BD)[27]. LDE and BD are calculated for the $4$ pairs of points($A-I$, $B-H$, $C-G$ and $D-F$) around $E$.
@@
In (5), $V=[A,B,C,D,F,G,H,I]$ is the vector of $8$ SST values around $E$. $V_{max}$ is the maximum of these components in $V$, $V_{min}$ is the minimum and $\overline{V}$ the average. It can be verified that LDE and BD are two mathematical operators which fall into $[0,1]$ for normalization[26, 25].
@@ II-E Skeletonization
$\begin{gathered}S(Z)=\bigcup_{k=0}^{K}S_{k}(Z)\\ S_{k}(Z)=(Z\ominus kM)-(Z\ominus k)\degree M\\ K=\max\{k|(Z\ominus kM)\neq\varnothing\}\end{gathered}$ (10)
The principle can be seen in figure 3, and the mathematical principle can be seen in (10). In it, $Z$ is the determined frontal zone, $(Z\ominus kM)$ means performing $k$ rounds of corrosion on the identified frontal zone, $M$ is the convolutional kernel, °$\degree$ is the open operation and $K$ is the maximum number of corrosion cycles before being corroded into an empty set.
@@ II-F Skeleton trimming
Although these fronts have been found, they have many branches, and in order to extract the main features and adapt to the research scale, redundant branches must be removed. Here, a greedy algorithm called the discrete skeleton evolution(DSE) algorithm is used to assign weights to each branch, iteratively removing the branch with the minimum weight to obtain the main fronts[28, 29].
@@ IV-B1 Detection
The higher occurrence of SST frontal zones in the northern parts of SCS, particularly in the latitudes between $10\degree$N and $20\degree$N, reflects seasonal transitions, where the differences between warm and cold ocean currents generate frontal zones. The seasonal shift in atmospheric circulation, coupled with oceanic upwelling and the influence of wind patterns, plays a crucial role in the formation of these frontal zones. The convergence of different water masses, such as warm tropical waters and cold subpolar currents, is responsible for the intensification of these frontal zones.
@@
Winter shows the highest intensity ($0.25\degree$C/km), followed by spring ($0.20\degree$C/km) summer ($0.15\degree$C/km) and fall ($0.14\degree$C/km). Intensity refers to the temperature gradient across the ocean front, which is typically stronger in spring and winter. Spring has more distinct temperature differences due to seasonal transitions, while winter’s cold temperatures can create higher gradients. The lower intensity values in summer and fall indicate less temperature contrast between the ocean and surrounding atmosphere.
@@ IV-C Comparison
TABLE II: QUANTITATIVE COMPARISONS ON SST DATASET
@@
Table II presents a quantitative comparison of the four methods, the value is the average of detection results from 2022 to 2024. As to number, improved method significantly outperforms the other methods, detecting only $71.44$ fronts, which indicates a more accurate and less over-detecting performance. In contrast, histogram method detects $263.22$ fronts, indicating that it tends to over-detect front, leading to potential false results. Lyapunov and gradient method also detect more fronts than improved method, but still show better accuracy compared to histogram method.
Improved method yields the highest intensity of $0.16\degree$C/km, indicating a sharper detection of the fronts. Gradient method comes close with $0.093\degree$C/km, but both histogram and Lyapunov method have much lower intensity values, suggesting weaker and less defined front detection. This suggests that improved method is more capable of detecting stronger fronts, while histogram and Lyapunov methods tend to miss or under-represent the intensity of these fronts.
@@ V Conclusion
Quantitative comparisons demonstrate the superiority of BFDT-MSA: it reduces over-detection by $73\%$ compared to histogram-based approaches, achieves sharper intensity ($0.16\degree$C/km), and captures broader ($40.56$ km) and longer ($212.42$km) fronts. These improvements stem from the fusion of multi-scale information and iterative refinement strategies, which traditional gradient or Lagrangian methods lack. Additionally, the open-source release of the tracking algorithm fills a critical gap in reproducible research.